RC meeting and Run Plan, Saturday January 21st
Return to the RC Meeting Calendar
RC Meeting Information
|Current RC:||Iuliia Skorodumina|
|Time:||No RC meeting today|
Friday Swing (01/20). Shift started in Controlled Access as the accelerator was still recovering from the repair and the HCal work was still ongoing in the hall. Around 6pm we requested Beam Permit and SBS on. MCC had troubles with turning the Correctors on, so we went back to Controlled Access and Jack Segal went down to the hall to unlock them. By 7:30pm we were in Beam Permit with all the magnets ON and waiting for the beam.
The accelerator recovery took longer than expected, so there was no beam until the end of Swing Shift. Swing Shift Summary (https://logbooks.jlab.org/entry/4113894).
Saturday Owl (01/21). At around 2am MCC started sending us CW beam and we started the beam checkout. First harp scans were not very good, but with the optics adjustments from MCC soon we got a good beam spot on target (0.46 and 0.48, which is within our new specifications, see https://logbooks.jlab.org/entry/4114126). We proceeded with the Carbon Hole run and checked the raster. The raster looked ok, so around 5am we went to production. Then we had a few complications.
First, there was a DAQ issue (https://logbooks.jlab.org/entry/4114174 and https://logbooks.jlab.org/entry/4114187). Thanks to the detailed How-to instructions, the Shift Crew was able to resolve the problem.
Then, we had a problem with the target control (https://logbooks.jlab.org/entry/4114223 and https://logbooks.jlab.org/entry/4114227). It turned out that iocha26 crashed (https://logbooks.jlab.org/entry/4114228). Thanks to the detailed instructions from Arun, we managed to reboot it (https://logbooks.jlab.org/entry/4114254 and https://logbooks.jlab.org/entry/4114256).
Despite the difficulties, we were able to run production overnight. We took a few runs at 20 uA and at 35 uA and then went up to 45 uA around 8am. Owl Shift Summary (https://logbooks.jlab.org/entry/4114044). We plan to continue production at 45 uA during the Day Shift.
Saturday Day (01/21). Taking production data the whole shift. Started at 45 uA, but at 10:30 am went down to 40 uA for the sake of better beam stability. The current plan is to stay at 40 uA for the whole weekend.
See the logentry from Ralph with detailed instructions on troubleshooting EPICs GUI malfunction (https://logbooks.jlab.org/entry/4114374). The How-To wiki instructions have been updated accordingly.
See also Charge Asymmetry Measurements from Paul (https://logbooks.jlab.org/entry/4114309).
Day Shift summary (https://logbooks.jlab.org/entry/4114427).
- Hodoscope issue reappeared (https://logbooks.jlab.org/entry/4114198 and https://logbooks.jlab.org/entry/4114201). This was an EPICs software issue and Mark already fixed it (https://logbooks.jlab.org/entry/4114303 and https://logbooks.jlab.org/entry/4114331).
- GEM Layer 2 efficiency is low, but this layer probably needs the software alignment after yesterday's work in the hall (https://logbooks.jlab.org/entry/4114218). Addressed -> the GEMs software alignment has been performed and the GEMs plots now look good (see Sean's logentry https://logbooks.jlab.org/entry/4114372).
- Some issue with the BBCal threshold readback (https://logbooks.jlab.org/entry/4114196), which is fixed already (https://logbooks.jlab.org/entry/4114356).
- Continue production on pol He3 cell at 40 uA.
We need Shift Leaders for the next three Owl Shifts. Please consider signing up .